# 2d Ising Critical Couplings from Quantum Gravity

### v1

@@ Abstract
Using an exact holographic duality formula between the inhomogeneous 2d Ising model and 3d quantum gravity, we provide a formula for ”real” zeroes of the 2d Ising partition function on finite graphs in terms of the geometry of a 2d triangulation embedded in the three-dimensional Euclidean space. The complex phase of those zeroes is given by the dihedral angles of the triangulation, which reflect its extrinsic curvature within the ambient 3d space, while the modulus is given by the angles within the 2d triangles, thus encoding the intrinsic geometry of the triangulation. Our formula can not cover the whole set of Ising zeroes, but we conjecture that a suitable complexification of these ”real” zeroes would provide a more thorough formula. Nevertheless, in the thermodynamic limit, in the case of flat planar 2d triangulations, our Ising zeroes formula gives the critical couplings for isoradial graphs, confirming its generality. This approach shows an intricate, but precise, new relation between statistical mechanics and quantum geometry.
@@
The paper is organized as follows. We first review the duality formula Bonzom:2015ova . We recall the basics of the 2d Ising model and of boundary spin network states of quantum gravity. We introduce a specific class of coherent states, such that the quantum gravity amplitudes are the inverse square of the 2d Ising partition function. Then a second section is dedicated to deriving a formula for zeroes of the Ising partition function from the semi-classical regime of the Ponzano-Regge state-sum and providing them with a parametrization in terms of 2d triangulations. We derive, what we refer to, as ”real” zeroes and conjecture that complexifying our formula should provide a more general ansatz. We conclude with a comparaison with known results on Ising zeroes and critical couplings, and discuss possible extensions of the present work.
@@ III Zeroes & Critical Couplings
To summarize, we have used the BCL duality between 3d quantum gravity and the 2d inhomogeneous Ising model, put in light in Bonzom:2015ova and based on Westbury formula Westbury1998 , to design a new method to find the zeroes of the Ising model. This duality relates the Ponzano-Regge path integral amplitudes for 3d quantum gravity for a bounded region of space-time to the 2d Ising model on the boundary surface, where the Ising couplings are given in terms of the coherent state parameters determining the quantum state of geometry on the boundary. More precisely, the Ponzano-Regge state-sum is equal, up to factors, to the inverse square of the Ising partition function. This allowed us to describe the 2d Ising zeroes in terms of the semi-classical regime of quantum gravity, using asymptotic saddle points of the Ponzano-Regge amplitudes at large spins. This led us to a new formula (23) for the Ising zeroes. Remarkably, it is consistent with the formula for Ising’s critical couplings on isoradial graphs and it extends the latter into an expression for the critical couplings in terms of the geometry of 2d triangulations embedded in the flat 3d space.

### v2

@@ Abstract
Using an exact holographic duality formula between the inhomogeneous 2d Ising model and 3d quantum gravity, we provide a formula for ”real” zeroes of the 2d Ising partition function on finite trivalent graphs in terms of the geometry of a 2d triangulation embedded in the three-dimensional Euclidean space. The complex phase of those zeroes is given by the dihedral angles of the triangulation, which reflect its extrinsic curvature within the ambient 3d space, while the modulus is given by the angles within the 2d triangles, thus encoding the intrinsic geometry of the triangulation. Our formula can not cover the whole set of Ising zeroes, but we conjecture that a suitable complexification of these ”real” zeroes would provide a more thorough formula. Nevertheless, in the thermodynamic limit, in the case of flat planar 2d triangulations, our Ising zeros’ formula gives the critical couplings for isoradial graphs, confirming its generality. Finally, the formula naturally extends to graphs with arbitrary valence in terms of geometry of circle patterns embedded in 3d space. This approach shows an intricate, but precise, new relation between statistical mechanics and quantum geometry.
@@
The paper is organized as follows. We first review the duality formula Bonzom:2015ova . We recall the basics of the 2d Ising model and of boundary spin network states of quantum gravity. We introduce a specific class of coherent states, such that the quantum gravity amplitudes are the inverse square of the 2d Ising partition function. Then a second section is dedicated to deriving a formula for zeros of the Ising partition function from the semi-classical regime of the Ponzano-Regge state-sum and providing them with a parametrization in terms of 2d triangulations. We derive, what we refer to, as ”real” zeroes and conjecture that complexifying our formula should provide a more general ansatz. We further extend our analysis from 3-valent graphs, dual to 2d triangulations, to graphs with nodes of arbitrary higher valence, in which case our geometric formula for Ising zeros is written in terms of circle patterns embedded in 3d space. We conclude with a comparaison with known results on Ising zeroes and critical couplings, and discuss possible extensions of the present work.
@@ III Zeroes & Critical Couplings
IV Higher Valence and Circle Patterns
We would like to conclude our analysis by showing that, assuming that our formula is indeed correct for 3-valent planar graphs, then it straightforwardly extends to planar graphs with arbitrary valence, in terms of the geometry of circle patterns embedded in the flat 3d space.
Indeed, while a 3-valent node is naturally dual to a triangle, a node with higher valence is dual to a polygon with as many edges as links connected to the node. Then a planar graph with nodes of arbitrary valence (higher or equal to 3) can be embedded in 3d space as dual to a 2d cellular complex with the topology of a 2-sphere, as illustrated on fig.9.
[…]
Figure 9: 2d cellular complex, with the topology of a 2-sphere, made of polygons glued together, and its dual graph in blue.
A polygon can be decomposed into triangles by choosing one of its vertices and drawing lines from it to all its other vertices. As drawn on fig.10, this unfolds the corresponding graph node into a tree of 3-valent nodes. We will refer to the new links between these 3-valent nodes as internal links. Doing so at every node turns the original graph $\Gamma$ into a planar 3-valent graph $\overline{\Gamma}$, to which we can now apply our Ising zeros’ formula. The key point is that the loop polynomial on the original graph is (obviously) exactly equal to the the loop polynomial on the unfolded 3-valent graph, if setting all the Ising coupling $Y_{\ell}$ on the internal links to 1:
$P_{\Gamma}[\{Y_{\ell}\}_{\ell\in\Gamma}]=P_{\overline{\Gamma}}[\{Y_{\ell}\}_{\ell\in\Gamma},\{Y_{\ell}=1\}_{\ell\textrm{ internal}}]\,.$ (27)
Figure 10: Decomposition of a polygon into triangles and its dual unfolding of a higher valent node into a tree of 3-valent nodes. Internal triangulation edges ad their dual internal links are in dotted black and dotted blue.
Therefore we get zeros of the loop polynomial on the graph $\Gamma$ from the loop polynomial on the extended graph $\overline{\Gamma}$. The latter has only 3-valent nodes and we can apply our geometric ansatz (23) for Ising zeros. We thus need to understand the geometrical meaning of setting $Y_{\ell}=1$ on a link. Looking at the formula (23), one sees that the coupling $Y_{\ell}$ is equal to 1 if and only if its phase given by the dihedral angle vanishes, $\theta_{\ell}=0$, and the product of the tangent of the half triangle angle is equal to 1, i.e. if the sum of the opposite triangle angles is equal to $\varphi_{\ell}+\widetilde{\varphi}_{\ell}=\pi$. This is automatically the case for two neighbouring triangles if the polygon is inscribed in a circle, as drawn on fig.11. This provides a natural geometric (sufficient) condition for gluing 3-valent Ising vertices into a higher valent vertex.
Figure 11: Two neighbouring triangles in a polygon inscribed in circle. The geometric coupling $Y_{\ell}$ on the internal link $\ell$ between the two triangles is automatically equal to 1, as the dihedral angle vanishes $\theta_{\ell}=0$ and the opposite triangle angles sum to $\varphi_{\ell}+\widetilde{\varphi}_{\ell}=\pi$.
So let us assume that every polygon is inscribed in a circle, as drawn on fig.11. This is called a circle pattern. Work on Ising critical couplings on circle patterns was done e.g. in lis2019circle (see also chelkak2022ising ). The extra ingredient of the present analysis is that the circle pattern is not drawn in the flat 2d space but embedded in the flat 3d space, and that the dihedral angles between the polygon planes play the key role of determining the phases of the Ising couplings. Indeed, our formula for the Ising zeros of the 3-valent graphs leads directly to a generalized formula on the graph $\Gamma$ with arbitrary valence. The ansatz for the couplings on the links of the graph (and not on the internal link, which have been set to 1) is:
$\forall\ell\in\Delta\,,\,\,Y_{\ell}^{(0)}=e^{i\epsilon\frac{\theta_{\ell}}{2}}\,\sqrt{\tan\frac{\psi_{\ell}}{4}\tan\frac{{\widetilde{\psi}_{\ell}}}{4}}\,,$ (28)
where $\psi_{\ell}$ and $\tilde{\psi}_{\ell}$ are now the angles at the center of the two circles sharing the edge dual to the link $\ell$, as drawn on fig. 12. Due to basic trigonometry, these are automatically equal to the opposite triangle angle whatever the chosen decomposition of the polygon into triangles. The angle $\theta_{\ell}$ remains the dihedral angle between the normal vectors to the two polygons linked by $\ell$ (or equivalently sharing its dual edge).
Figure 12: Triangle angles and center angles for the Ising zeros’ geometric ansatz (28) on embedded circle patterns. Whatever the decomposition of the polygon into triangles, the center angle $\psi_{\ell}$ for a graph link $\ell$ (not an internal link) is always twice the opposite triangle angle $\varphi_{\ell}$. Thus the generalized formula (28) for higher valent graphs naturally follows from the geometric ansatz (23) for 3-valent graphs and their dual 2d triangulations.
This shows that the geometric formula for Ising zeros from quantum gravity naturally extends from 3-valent graphs (dual to triangulations) to graphs with nodes of arbitrary higher valence embedded in circle patterns, but also underlines the potentially interesting role of circle patterns on the boundary of the Ponzano-Regge path integral for 3d quantum gravity.
@@ IV Higher Valence and Circle Patterns
To summarize, we have used the BCL duality between 3d quantum gravity and the 2d inhomogeneous Ising model, put in light in Bonzom:2015ova and based on Westbury formula Westbury1998 , to design a new method to find the zeros of the Ising model. This duality relates the Ponzano-Regge path integral amplitudes for 3d quantum gravity for a bounded region of space-time to the 2d Ising model on the boundary surface, where the Ising couplings are given in terms of the coherent state parameters determining the quantum state of geometry on the boundary. More precisely, the Ponzano-Regge state-sum is equal, up to factors, to the inverse square of the Ising partition function. This allowed us to describe the 2d Ising zeroes in terms of the semi-classical regime of quantum gravity, using asymptotic saddle points of the Ponzano-Regge amplitudes at large spins. This led us to a new formula (23) for the Ising zeroes. Remarkably, it is consistent with the formula for Ising’s critical couplings on isoradial graphs and it extends the latter into an expression for the critical couplings in terms of the geometry of 2d triangulations embedded in the flat 3d space. Moreover, we were able to extend our geometric ansatz from 3-valent graphs, dual to 2d triangulations, to graphs with arbitrary valence, dual to general 2d circle patterns.
@@
Acknowledgement
The authors would like to thank Marcin Lis for pointing out works studying the Ising model on circle patterns.